40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 11971     Entry time: Tue Feb 2 18:54:02 2016     In reply to: 11970     Reply to this: 11973
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: Lightwave NPRO moved from PSL table to SP table 

jiIn fact, it is one of the most difficult type mode profiling to measure a beam directly out from a laser source.

If you reduce the power by ADJ, this significantly changes the output mode as the pumping power varies temperature gradient of the laser crystal and thus thermal lensing in it. I'd recommend you to keep the nominal power.

If you use a PBS for power reduction, you should increase the transmission ~x10 from the minimum so that you are not dominated by possible junk polarization.

Any transmissive BK7 components where the beam is small can cause thermal lensing. In order to avoid this issue, I usually use two noncoated (or one AR coated) optical windows made of UV fused silica to pick off the beam. Once the beam power is reduced I suppose it is OK to use an additional ND filter in front of the CCD.

Another more reliable method is an old-good knife edge measurement.

ELOG V3.1.3-